OpenStudy (anonymous):

last check for the night... x^6 + x^2 _______________ x x^6+x ? @satellite73 and @jim_thompson5910

geerky42 (geerky42):

Factor x out in numerator. \[\dfrac{x^6 + x^2}{x} = \dfrac{x(x^5+x)}{x} = \boxed{x^5+x}\] Does this make sense?

OpenStudy (anonymous):

no, how did you get 5?

geerky42 (geerky42):

Well, I factored x out. \[x^6 = x\cdot x^5\]

geerky42 (geerky42):

\[\large x = x^1\]\[\large x\cdot x^5 = x^{5+1} = x^6\]

jimthompson5910 (jim_thompson5910):

or you can divide x^6 by x to get x^5 then you can divide x^2 by x to get x so that's another way to get the answer of x^5 + x

OpenStudy (anonymous):

Ok... i see where I went wrong... thank you both. :)

geerky42 (geerky42):

Yep. \[\dfrac{x^6 + x^2}{x} = \dfrac{x^6}{x}+\dfrac{x^2}{x}\]

geerky42 (geerky42):

Glad we helped.